OpenStudy (anonymous):

Fast limit:

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0^{-}}\frac{ e^{\frac{ 1 }{ x }} }{ x^2 }\]

OpenStudy (arabpride):

I have no idea o.o' Sorry :/ Wish I c00d help.

OpenStudy (anonymous):

I have an idea but I'm stuck: \[L'HOPITAL:\left[ \frac{ 0 }{ 0} \right]=>\lim_{x \rightarrow 0^{-}}\frac{ \frac{ -e^\frac{ 1 }{ x } }{ x^2 } }{ 2x }\] It's getting worse to a nonsense!

OpenStudy (anonymous):

@AravindG

OpenStudy (aravindg):

How did you get 0/0 form? e^infinity=infinity

OpenStudy (anonymous):

e^(-∞)=0 ln(0)=-∞

OpenStudy (anonymous):

Looks like nobody's into calculus on this platform

OpenStudy (anonymous):

No, there are people into calculus here.

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

\[ \frac{e^{1/x}}{x^2}= \exp\left(\ln\left(\frac{e^{1/x}}{x^2}\right)\right) = \exp\left(\frac 1x-2\ln(x)\right) \]

OpenStudy (anonymous):

This holds due to \(x>0\)

OpenStudy (anonymous):

x->0(-)=>x<0; For x>0 case the limit is simple is +∞;

OpenStudy (anonymous):

Wait, what I should have said is that \[ \frac{e^{1/x}}{x^2}>0 \]

OpenStudy (anonymous):

And if you are worried about the sign, then keep it as: \[ \ln(x^2) \]

OpenStudy (anonymous):

Or make it \(2\ln|x|\)

OpenStudy (anonymous):

wio still not helping we're at limits

OpenStudy (nincompoop):

ya it is

OpenStudy (anonymous):

\[ \lim_{x\to0^-}\exp\left(\frac 1x - 2\ln|x|\right) \]

OpenStudy (anonymous):

Finally I got it thanks wio